OpenStudy (anonymous):

how do you graph this system of equations x-y=3 x-2y=3

OpenStudy (anonymous):

Easy you first find what y and x equal by using elimination

OpenStudy (anonymous):

Do you know how to do elimination? (:

OpenStudy (anonymous):

yes, i think..

OpenStudy (anonymous):

You need me to clarify? (:

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

Haha yea sure (: Okay so you have to firsy see if anything can cancel out

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

As you can see nothing cancels out. Since your graphing it wont matter which one you start with. So then you have to think what do i have to multiply so that either the x or y variables will be crossed out

OpenStudy (anonymous):

I would try to cross the Y variable out first. So you Have to Multiply the whole equation (x - y = 3) by 2 so that the y will cross out so then

OpenStudy (anonymous):

I take that back you would multiply by -2

OpenStudy (anonymous):

So that you would be able to cross them out ( negative & positive)

OpenStudy (anonymous):

Okay so then you would get 2x + 2y = 6

OpenStudy (anonymous):

If you set them up like so 2x + 2y = 6 x - 2y = 3 You would see that the 2y does no indeed, cancel out. So then you'd get

OpenStudy (anonymous):

does* not does not lol

OpenStudy (anonymous):

So then if you add the two equations you get 3x = 9 x = 3 Then to solve for Y You would have

OpenStudy (anonymous):

\[\text{Psst, 'graph' $\not=$ 'solve'} \]

OpenStudy (anonymous):

Haha i know. But this is also a good way to do it. Okay so then you would plug the 3 in for x

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

No offence, but this is a very long way to do it.

OpenStudy (anonymous):

Its the way i was taught....

OpenStudy (anonymous):

Why dont YOU help him then :P

OpenStudy (anonymous):

i really appreciate the help

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is y= -3?

OpenStudy (anonymous):

Yes, im pretty sure (: Go ahead and show me how you got that and i will tell you if its correct. Btw sorry Newton I just have an immense respect for my teachers ^_^ Even though most the time I dont like them

OpenStudy (anonymous):

you put 3 in for x in one of the equations?

OpenStudy (anonymous):

Your teachers taught you that is how to figure out how to graph an equation?

OpenStudy (anonymous):

No they taught me to find the variables and then graph... lol easy peasy..

OpenStudy (anonymous):

The steps just seem long, after you did it a lot its not that hard :P

OpenStudy (anonymous):

do*

OpenStudy (toxicsugar22):

hi can anyone here help me after u help her

OpenStudy (anonymous):

Yes dark lol.. (: Okay and then you do y-intercept and slope to graph (:

OpenStudy (anonymous):

\[x-y = 3 \implies y = x-3 \] \[x-2y = 3 \implies y = \frac{x-3}{2} \] You (should be able to!) can graph these right form here, as they are in the form y = mx + c. But setting x = 0, y = 0 to find the intercepts may help.

OpenStudy (anonymous):

im a guy* thnak you! lol

OpenStudy (anonymous):

x - slope and y = y intercept

OpenStudy (anonymous):

Ik your a guy.. haha

OpenStudy (toxicsugar22):

oh srry can someone help me after him

OpenStudy (anonymous):

toxicsugar said her*........ anyways.................thank you very much!

OpenStudy (anonymous):

Haha oh.. ^_^ blonde moment... And yes good process. That's just the way i was taught ^_^